40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 12 10:57:38 2009, rob, Update, General, some stuff 
    Reply  Mon Jan 12 18:53:03 2009, Yoichi, Update, LSC, AS CCD centering and ASDD demod phase 
       Reply  Tue Jan 13 18:59:09 2009, Kakeru, Update, LSC, AS CCD centering and ASDD demod phase 
          Reply  Wed Jan 21 16:18:17 2009, Kakeru, Update, LSC, AS CCD centering and ASDD demod phase 
             Reply  Wed Jan 21 22:53:08 2009, rana, Update, LSC, AS CCD centering and ASDD demod phase 
Message ID: 1222     Entry time: Mon Jan 12 10:57:38 2009     Reply to this: 1223
Author: rob 
Type: Update 
Category: General 
Subject: some stuff 

The AS beam was not hitting the AS166 diode, so I aligned the last little steering mirror and adjusted the phase for MICH locking.

I turned on the HV supplies for the OMC.

Then I realigned the beam onto the AS166 diode, since the steering mirrors came on when I turned on the HV supplies.

It took awhile to find the alignment of the beam into the OMC. Once that was done, the output beam alignment was set, so I aligned onto the AS166 diode a third time.

The bottom two Sorensens in the OMC voltage supply don't look right. They have stickers that say +-24V, but each is sitting at 17.5V and showing no current draw. What's going on here?
ELOG V3.1.3-